Solve this proportion:
Solve for X:
x-3/x = 4/7

Answer by checkley75(3666)

x-3/x=4/7
(x^2-3)/x=4/7
7x^2-21=4x
7x^2-4x-21=0
using the quadratic equation we get:
x=(4+-sqrt[-4^2-4*7*-21])/2*7
x=(4+-sqrt[16+588])/14
x=(4+-sqrt604)/14
x=(4+-24.5764)/14
x=(4+24.5764)/14
x=28.5764/14
x=2.04 answer.
x=(4-24.5764)/14
x=-1.47 answer.